# ValuesRAG: Enhancing Cultural Alignment Through Retrieval-Augmented Contextual Learning

### v1

@@ Abstract
Cultural values alignment in Large Language Models (LLMs) is a critical challenge due to their tendency to embed Western-centric biases from training data, leading to misrepresentations and fairness issues in cross-cultural contexts. Recent approaches, such as role-assignment and few-shot learning, often struggle with reliable cultural alignment as they heavily rely on pre-trained knowledge, lack scalability, and fail to capture nuanced cultural values effectively. To address these issues, we propose ValuesRAG, a novel and effective framework that applies Retrieval-Augmented Generation (RAG) with in-context learning to integrate cultural and demographic knowledge dynamically during text generation. Leveraging the World Values Survey (WVS) dataset, ValuesRAG first generates summaries of values for each individual. Subsequently, we curated several representative regional datasets to serve as test datasets and retrieve relevant summaries of values based on demographic features, followed by a reranking step to select the top-k relevant summaries. ValuesRAG consistently outperforms baseline methods, both in the main experiment and in the ablation study where only the values summary was provided, highlighting ValuesRAG’s potential to foster culturally aligned AI systems and enhance the inclusivity of AI-driven applications.
@@ 1 Introduction
The rapid advancement of Large Language Models (LLMs) has revealed pressing challenges in cultural values alignment  Singh et al. (2024); Kharchenko et al. (2024); Hu et al. (2024). Predominantly trained on Western data sources Achiam et al. (2023); Touvron et al. (2023); Jiang et al. (2023), LLMs inherently reflect Western cultural norms and social biases, raising concerns about their applicability in global contexts. These biases present significant challenges when deploying LLMs in cross-cultural environments, often resulting in misrepresentations and stereotypical outputs Gallegos et al. (2024); Xie et al. (2024); Potter et al. (2024); Huang et al. (2024). Despite ongoing efforts to address these issues, existing strategies often fall short. While some countries have developed localized LLMs, such as China’s ERNIE Sun et al. (2021), South Korea’s HyperCLOVA Yoo et al. (2024), and the multilingual ChatGLM GLM et al. (2024), these models also exhibit biases inherited from their respective training datasets. As a result, cultural and social biases embedded in LLMs remain a critical concern, compelling researchers to explore more robust frameworks for cultural alignment Gallegos et al. (2024); Xie et al. (2024); Potter et al. (2024).
@@
Figure 1: Overview of different approaches for cultural alignment. Comparing two baseline methods, namely Role Assignment and Few-Shot Learning, and the proposed ValuesRAG framework.
@@
Recent studies have proposed several approaches, such as role-assignment approaches and few-shot learning techniques  Li et al. (2024a); AlKhamissi et al. (2024), to mitigate these cultural biases. However, these methods still face several challenges: (1) Role-assignment approaches, relying solely on the model’s pre-trained knowledge, provide pre-defined demographic information but fail to incorporate explicit values alignment text, which subsequently introduces stereotypes and biases rooted in Western-centric training data; (2) While offering example-based guidance, few-shot learning methods struggle to comprehensively capture the complex cultural values due to the limited correlation between different values dimensions, thus remain ineffective on values-related tasks that differ significantly from the examples; (3) In addition, these methods can only align with the values of a single individual, and singular values cannot represent the universal values of individuals with similar characteristics.
@@
To address these challenges, we propose ValuesRAG, a novel framework that utilizes Retrieval-Augmented Generation (RAG) and in-context learning to dynamically incorporate cultural knowledge during text generation (see Figure 1). Our framework leverages the World Values Survey (WVS) dataset Haerpfer et al. (2022), a globally recognized and comprehensive dataset that explores values across countries using rigorous social science methodologies. Specifically, we first generate summaries for each topic, followed by generating individuals’ summaries of values and demographic profiles in parallel. After constructing the knowledge base, we retrieve the top $100$ relevant summaries based on demographic features, followed by a reranking step to ensure the most relevant top-k summaries are selected. Finally, we utilize a reasoning LLM that filters the most relevant demographic profiles and applies step-by-step reasoning grounded in the retrieved values to generate the final answer to the question.
@@
We evaluate the performance of ValuesRAG by comparing it against several baseline approaches, including: (1) zero-shot inference, (2) role-assignment-only method Tao et al. (2024), (3) few-shot learning Choenni and Shutova (2024), and (4) a hybrid method combining (1) and (2). To ensure a comprehensive evaluation, we curated diverse regional survey QA datasets which are designed to capture values-related question-answer pairs. Extensive experimental results show significant improvements in cultural and contextual understanding, demonstrating that ValuesRAG outperforms the baselines, also in the ablation study where only summaries of values were provided without any additional information. Unlike previous methods that heavily depend on pre-trained knowledge or limited demonstrations, ValuesRAG dynamically retrieves and integrates multiple similar individual summaries based on demographic features, enabling richer value representations and more context-aware responses compared to approaches relying on a single predefined prompt or role.
@@
These findings highlight ValuesRAG’s potential to foster inclusive AI systems, enhancing the reliability and fairness of AI-driven applications. Our study demonstrates ValuesRAG’s robust capabilities on a global scale, suggesting its applicability in aligning the values of diverse groups within a single country. ValuesRAG provides a cost-efficient tool for public policymakers and scientists from various disciplines to refine social simulations, enabling more precise predictions of policy outcomes Li et al. (2024b). This, in turn, facilitates the creation of fairer and more effective policies. Moreover, NGOs can leverage ValuesRAG to develop LLMs that reflect specific value orientations while maintaining strong alignment with users’ values, thereby increasing their persuasive impact. This approach benefits the promotion and spread of values that contribute to the planet’s sustainable development and the long-term well-being of human society.
@@ 2.1 Evaluation of LLMs’ Cultural Bias
Pre-trained models are facing growing criticism for their inherent social biases, with cultural bias emerging as a particularly nuanced and pervasive issue Tao et al. (2024). Unlike the more obvious safety concerns and social discrimination Liu et al. (2024) embedded in language models, cultural bias manifests in subtler ways, often reflecting the dominant cultural perspectives present in training data. Studies have shown that LLMs often exhibit cultural biases aligned with the values of developed countries, resulting in the under-representation of perspectives from less developed regions  Manvi et al. (2024); Durmus et al. (2024). This imbalance not only perpetuates existing cultural hierarchies but also limits the global applicability of these models  Manvi et al. (2024). Various benchmarks and evaluation methods have been proposed to assess the cultural biases of pre-trained models Gallegos et al. (2024). For example, Webster et al.  Webster et al. (2021) developed probability-based metrics to evaluate gender bias embedded in pre-trained models, while Caliskan et al. Caliskan et al. (2017) pioneered the use of word embeddings as quantitative measures of bias. More recently, Karinshak et al. Karinshak et al. (2024) introduced LLM-GLOBE, a benchmark where LLMs generate both quantitative and open-ended answers to values assessment questions, with subsequent evaluation using the LLM-as-a-Jury Protocol. These evaluation methods collectively highlight the complex nature of cultural bias in LLMs and the need for multifaceted assessment approaches.
Set Dataset Abbreviation Region Year Na VQb
Train Set World Values Survey WVS Global 2017–2022 97.2k 259
Test Sets European Values Study EVS Europe 2017 59.4k 211
@@
Table 1: Overview of the datasets utilized in our study. The Train Set (WVS) includes global data collected between 2017 and 2022, providing the basis for generating cultural summaries of values and validation for our method. The Test Set consists of six region-specific surveys, each capturing socio-cultural information from distinct geographic areas and time frames. aN represents sample size in thousands (k). bVQ represents the number of values-related questions. cDQ represents the number of demographic features.
@@ 2.2 Mitigation of LLMs’ Cultural Bias
Methods like RLHF Shen et al. (2023); Ji et al. (2024) are very commonly used in LLM values alignment, but this type of single values alignment method is not suitable for the mitigation of cultural bias, because the alignment goal of cultural bias is diverse and dynamic Huang et al. (2024), as there are hundreds of countries and cultures in the earth. Addressing cultural biases has become a critical area of research, with various strategies being proposed to enhance cultural sensitivity in LLMs. For instance, Tao et al. Tao et al. (2024) adopted national and cultural role assignments to adjust the cultural values of LLMs, while Masoud et al. Masoud et al. (2024) developed a soft prompt tuning approach to mitigate bias. Also, Choenni and Shutova Choenni and Shutova (2024) employed few-shot in-context learning to align cultural behaviors, demonstrating promising results in specific contexts. However, these approaches face significant limitations in fully capturing the complexity of cultural alignment. Tao et al.’s technique mainly depends on national and cultural roles without explicitly integrating values assignments, causing an overreliance on latent internal representations. Meanwhile, Choenni and Shutova’s few-shot learning approach similarly falls short of modeling cultural alignment in all its complexity. We, therefore, use these methods as baselines to benchmark our proposed approach.
@@ 3 Datasets
In this section, we first introduce the World Values Survey (WVS) as our training dataset (Section 3.1), highlighting its extensive coverage, global representativeness, and relevance for values-related studies. Subsequently, we describe six regional test datasets (Section 3.2), which are carefully selected to ensure geographic, cultural, and demographic diversity.
3.1 Train Dataset
WVS  [^1] https://www.worldvaluessurvey.org/wvs.jsp Haerpfer et al. (2022) is a globally recognized dataset that investigates human beliefs, values, and cultural norms through structured surveys conducted across multiple countries. WVS is selected as our train dataset especially due to its numerous advantages:
@@ 1. Broad Recognition and Inclusiveness:
WVS is widely recognized and used by governments, social scientists, and international organizations in values studies. It covers 120 countries, representing 94.5% of the global population, ensuring broad geographic and cultural representation.
@@ 2. Expert-Designed and Accessible:
The dataset is designed by leading domain experts to conduct comprehensive surveys of values, ensuring reliability and relevance. It is publicly accessible, enabling reproducibility and transparency in research.
@@ 3. Effective Structure and Large Scale:
Since values evolve gradually over time, WVS is conducted in waves, with each wave occurring every five years. For our study, we utilized the most recent wave, spanning from 2017 to 2022.
The WVS codebook includes over 600 indicators, with 259 values-related and 31 demographic-related questions. The values questions span 13 topics, such as social trust, post-materialism, and political interest. We randomly selected 20% (52 questions) per topic for validation and used the remaining 80% (207 questions) for summary generation. The 31 demographic features, including country, sex, age, education, social class, and employment status, were used to generate demographic summaries for retrieval tasks.
@@ 3.2 Test Datasets
We selected six regional surveys to serve as test datasets based on the following criteria:
@@ 2. Temporal Proximity:
The datasets exhibit temporal proximity to WVS Wave 7 (2017–2022), allowing aligned comparisons and ensuring consistency across evaluations.
The regions in our test datasets were meticulously chosen to encompass a wide range of geographic, cultural, and demographic diversity, ensuring the data accurately reflects the majority of the global population. All of them are publicly accessible, and are statistically representative at national or regional levels, which guarantees their reliability and validity.
To be specific, we selected European Values Study EVS (2022) as the representative dataset for Europe, as it is the largest values survey in the region. For the United States, we selected the General Social Survey Davern et al. (2024), which is the most comprehensive social survey in the country. The Chinese General Social Survey Bian and Li (2012) serves as the representative dataset for China due to its comprehensive sampling methodology and scientific rigor. For India, where national survey data were largely inaccessible during our study, we used Pew Research Center’s survey data Sahgal and Evans (2021) to represent the Indian population. The AmericasBarometer Lab (2021), conducted by the LAPOP Lab, was selected to represent Latin American countries, as it covers 32 countries across the region. Finally, Afrobarometer Afrobarometer (2023) was chosen as the representative dataset for Africa. A detailed summary of these datasets is presented in Table 1.
Figure 2: Overview of the proposed ValuesRAG framework for cultural alignment, consisting of four components: (1) Dataset Preprocessing, (2) Topic QA Summary Generation, (3) Individual Values Summary Generation, and (4) Values Retrieval Augmented Generation.
@@ 4 Methodology
In this section, we present the ValuesRAG which is specifically designed to address cultural biases and enhance contextual alignment in LLM-driven scenarios through a Retrieval-Augmented Generation (RAG) approach. ValuesRAG consists of three key components: Values and Demographic Summary Generation, which extracts and summarizes cultural values and demographic information from large-scale datasets (Section 4.1); Values-Augmented Generation, which incorporates these summaries into the generative process to align responses with the cultural context (Section 4.2); and Retrieval-based Values Alignment, which dynamically assigns relevant individual values to queries based on demographic profiles (Section 4.3). An overview of the ValuesRAG framework is provided in Figure 2.

### v2

@@ Abstract
Cultural values alignment in Large Language Models (LLMs) is a critical challenge due to their tendency to embed Western-centric biases from training data, leading to misrepresentations and fairness issues in cross-cultural contexts. Recent approaches, such as role-assignment and few-shot learning, often struggle with reliable cultural alignment as they heavily rely on pre-trained knowledge, lack scalability, and fail to capture nuanced cultural values effectively. To address these issues, we propose ValuesRAG, a novel and effective framework that applies Retrieval-Augmented Generation (RAG) with In-Context Learning (ICL) to integrate cultural and demographic knowledge dynamically during text generation. Leveraging the World Values Survey (WVS) dataset, ValuesRAG first generates summaries of values for each individual. Subsequently, we curate several representative regional datasets to serve as test datasets and retrieve relevant summaries of values based on demographic features, followed by a reranking step to select the top-k relevant summaries. ValuesRAG consistently outperforms baseline methods, both in the main experiment and in the ablation study where only the values summary was provided. Notably, ValuesRAG demonstrates an accuracy of 21% improvement over other baseline methods, highlighting its potential to foster culturally aligned AI systems and enhance the inclusivity of AI-driven applications.
@@ 1 Introduction
The rapid advancement of Large Language Models (LLMs) has revealed pressing challenges in cultural values alignment  Singh et al. (2024); Kharchenko et al. (2024); Hu et al. (2024). Predominantly trained on Western data sources Achiam et al. (2023); Touvron et al. (2023); Jiang et al. (2023), LLMs inherently reflect Western cultural norms and social biases, raising concerns about their applicability in global contexts. These biases present significant challenges when deploying LLMs in cross-cultural environments, often resulting in misrepresentations and stereotypical outputs Gallegos et al. (2024); Xie et al. (2024); Potter et al. (2024); Huang et al. (2024). Despite ongoing efforts to address these issues, existing strategies often fall short. While some countries have developed localized LLMs, such as China’s ERNIE Sun et al. (2021), ChatGLM GLM et al. (2024), DeepSeek Liu et al. (2024a), and South Korea’s HyperCLOVA Yoo et al. (2024), these models also exhibit biases inherited from their respective training datasets. As a result, cultural and social biases embedded in LLMs remain a critical concern, compelling researchers to explore more robust frameworks for cultural alignment Gallegos et al. (2024); Xie et al. (2024); Potter et al. (2024).
@@
Figure 1: Overview of different approaches for cultural alignment. Comparing two baseline methods, namely Role Assignment and Few-Shot Learning, and our proposed ValuesRAG framework.
@@
Recent studies have proposed several approaches, such as role-assignment approaches Tao et al. (2024) and few-shot learning techniques  Choenni and Shutova (2024), to mitigate these cultural biases. However, these methods still face several challenges: (1) Role-assignment approaches, relying solely on the model’s pre-trained knowledge, provide pre-defined demographic information but fail to incorporate explicit values alignment text, which subsequently introduces stereotypes and biases rooted in Western-centric training data; (2) While offering example-based guidance, few-shot learning methods struggle to comprehensively capture the complex cultural values due to the limited correlation between different values dimensions, thus remain ineffective on values-related tasks that differ significantly from the examples; (3) In addition, these methods can only align with the values of a single individual, and singular values cannot represent the universal values of individuals with similar characteristics.
@@
To address these challenges, we propose ValuesRAG, a novel framework that utilizes Retrieval-Augmented Generation (RAG) and In-Context Learning (ICL) to dynamically incorporate cultural knowledge during text generation (see Figure 1). Our framework leverages the World Values Survey (WVS) dataset Haerpfer et al. (2022), a globally recognized and comprehensive dataset that explores values across countries using rigorous social science methodologies. Specifically, we first generate summaries for each topic, followed by generating individuals’ summaries of values and demographic profiles in parallel. After constructing the knowledge base, we retrieve the top $100$ relevant summaries based on demographic features, followed by a reranking step to ensure the most relevant top-k summaries are selected. Finally, we utilize a reasoning LLM that filters the most relevant demographic profiles and applies reasoning grounded in the retrieved values to generate the final answer to the question.
@@
We evaluate the performance of ValuesRAG by comparing it against several baseline approaches, including: (1) zero-shot inference, (2) role-assignment-only method Tao et al. (2024), (3) few-shot learning Choenni and Shutova (2024), and (4) a hybrid method combining (1) and (2). To ensure a comprehensive evaluation, we curated diverse regional survey QA datasets which are designed to capture values-related question-answer pairs. Extensive experimental results show significant improvements in cultural and contextual understanding, demonstrating that ValuesRAG outperforms the baselines. Unlike previous methods that heavily depend on pre-trained knowledge or limited demonstrations, ValuesRAG dynamically retrieves and integrates multiple similar individual values summaries based on demographic features, enabling richer value representations and more context-aware responses compared to approaches relying on a single predefined prompt or role.
In addition, ablation studies on using only value-augmented generation and on varying the number of retrieved summaries confirm ValuesRAG’s robust performance under different configurations. ValuesRAG surpasses the baselines through purely values-based generation. Meanwhile, adjusting the number of retrieved documents shows that moderate retrievals can effectively balance diversity and relevance, also maintaining high accuracy across multiple benchmarks.
@@
These findings highlight ValuesRAG’s potential to foster inclusive AI systems, enhancing the reliability and fairness of AI-driven applications. Our study demonstrates ValuesRAG’s robust capabilities on a global scale, also suggesting its applicability in aligning the values of diverse groups within a single country. ValuesRAG provides a cost-efficient tool for public policymakers and scientists from various disciplines to refine social simulations, enabling more precise predictions of policy outcomes Li et al. (2024). This, in turn, facilitates the creation of fairer and more effective policies. Moreover, NGOs can leverage ValuesRAG to develop LLMs that reflect specific value orientations while maintaining strong alignment with users’ values, thereby increasing their persuasive impact. This approach benefits the promotion and spread of values that contribute to the planet’s sustainable development and the long-term well-being of human society.
@@ 2.1 Evaluation of LLMs’ Cultural Bias
Pre-trained models are facing growing criticism for their inherent social biases, with cultural bias emerging as a particularly nuanced and pervasive issue Tao et al. (2024). Unlike the more obvious safety concerns and social discrimination Liu et al. (2024b) embedded in language models, cultural bias manifests in subtler ways, often reflecting the dominant cultural perspectives present in training data. Studies have shown that LLMs often exhibit cultural biases aligned with the values of developed countries, resulting in the under-representation of perspectives from less developed regions  Manvi et al. (2024); Durmus et al. (2024). This imbalance not only perpetuates existing cultural hierarchies but also limits the global applicability of these models  Manvi et al. (2024). Various benchmarks and evaluation methods have been proposed to assess the cultural biases of pre-trained models Gallegos et al. (2024). For example, Webster et al.  Webster et al. (2021) developed probability-based metrics to evaluate gender bias embedded in pre-trained models, while Caliskan et al. Caliskan et al. (2017) pioneered the use of word embeddings as quantitative measures of bias. More recently, Karinshak et al. Karinshak et al. (2024) introduced LLM-GLOBE, a benchmark where LLMs generate both quantitative and open-ended answers to values assessment questions, with subsequent evaluation using the LLM-as-a-Jury Protocol. These evaluation methods collectively highlight the complex nature of cultural bias in LLMs and the need for multifaceted assessment approaches.
Category Dataset Abbreviation Region Year N VQ
Retrieval Corpus World Values Survey WVS Global 2017–2022 97.2k 259
Test Datasets European Values Study EVS Europe 2017 59.4k 211
@@
Table 1: Overview of the datasets utilized in our study. The Retrieval Corpus (WVS) includes global data collected between 2017 and 2022, providing the basis for generating cultural summaries of values and validation for our method. The Test Datasets consist of six region-specific surveys, each capturing socio-cultural information from distinct geographic areas and time frames. N represents sample size in thousands (k). VQ represents the number of values-related questions.
@@ 2.2 Mitigation of LLMs’ Cultural Bias
Methods like RLHF Shen et al. (2023); Ji et al. (2024) are very commonly used in LLM values alignment, but this type of single values alignment method is not that suitable for the mitigation of cultural bias, because the alignment goal of cultural bias is diverse and dynamic Huang et al. (2024), as there are hundreds of countries and cultures in the earth. Addressing cultural biases has become a critical area of research, with various strategies being proposed to enhance cultural sensitivity in LLMs. For instance, Tao et al. Tao et al. (2024) adopted national and cultural role assignments to adjust the cultural values of LLMs, while Masoud et al. Masoud et al. (2024) developed a soft prompt tuning approach to mitigate bias. Also, Choenni and Shutova Choenni and Shutova (2024) employed few-shot in-context learning to align cultural behaviors, demonstrating promising results in specific contexts. However, these approaches face significant limitations in fully capturing the complexity of cultural alignment. Tao et al.’s technique mainly depends on national and cultural roles without explicitly integrating values assignments, causing an overreliance on latent internal representations. Meanwhile, Choenni and Shutova’s few-shot learning approach similarly falls short of modeling cultural alignment in all its complexity. We therefore use these methods as baselines to benchmark our proposed approach.
@@ 3 Datasets
In this section, we first introduce the World Values Survey (WVS) as our retrieval corpus (Section 3.1), highlighting its extensive coverage, global representativeness, and relevance for values-related studies. Subsequently, we describe six regional test datasets (Section 3.2), which are carefully selected to ensure geographic, cultural, and demographic diversity.
3.1 Retrieval Corpus
VS  [^1] https://www.worldvaluessurvey.org/wvs.jsp Haerpfer et al. (2022) is a globally recognized dataset that investigates human beliefs, values, and cultural norms through structured surveys conducted across multiple countries. WVS is selected as our retrieval corpus especially due to its numerous advantages:
@@ 1. Broad recognition and inclusiveness:
WVS is widely recognized and frequently used by governments, social scientists, and major international organizations in comparative values studies. It currently covers 120 countries, representing 94.5% of the global population, ensuring broad geographic and cultural representation.
@@ 2. Expert-designed and accessible:
The dataset is meticulously designed by leading domain experts to conduct comprehensive surveys of values, ensuring reliability, rigor, and relevance. It is publicly accessible, enabling reproducibility and transparency in research.
@@ 3. Effective structure and large scale:
Since values evolve gradually over time, VS is conducted in waves, with each wave occurring every five years. For our study, we utilize the most recent wave, spanning from 2017 to 2022.
The WVS codebook includes over 600 indicators, with 259 values-related and 31 demographic-related questions. The value questions span 13 topics, such as social trust, post-materialism, and political interest. We randomly select 20% (52 questions) per topic for validation and use the remaining 80% (207 questions) for summary generation. The 31 demographic features, including country, sex, age, education, social class, and employment status, are used to generate demographic summaries for retrieval tasks.
Figure 2: Overview of the proposed ValuesRAG framework for cultural alignment. Consisting four components: (1) Dataset Preprocessing, (2) Topic QA Summary Generation, (3) Individual Values Summary Generation, and (4) Values Retrieval Augmented Generation.
@@ 3.2 Test Datasets
We select six regional surveys [^2] Detailed dataset description are provided in Appendix A. to serve as test datasets based on the following criteria:
@@ 2. Temporal proximity:
The datasets exhibit close temporal proximity to WVS Wave 7 (2017–2022), thereby allowing aligned comparisons and ensuring thorough consistency across diverse global evaluations.
The regions in our test datasets are meticulously chosen to encompass a wide range of geographic, cultural, and demographic diversity, ensuring that the data accurately reflects the majority of the global population. All of them are publicly accessible and are statistically representative at national or regional levels, which guarantees their reliability and validity.
To be specific, we select the European Values Study EVS (2022) as the representative dataset for Europe, as it is the largest values survey in the region. For the United States, we select the General Social Survey Davern et al. (2024), which is the most comprehensive social survey in the country. The Chinese General Social Survey Bian and Li (2012) serves as the representative dataset for China due to its comprehensive sampling methodology and scientific rigor. For India, where national survey data were largely inaccessible during our study, we use Pew Research Center’s survey data Sahgal and Evans (2021) to represent the Indian population. The AmericasBarometer Lab (2021), conducted by the LAPOP Lab, is selected to represent Latin American countries, as it covers 32 countries across the region. Finally, Afrobarometer Afrobarometer (2023) is chosen as the representative dataset for Africa. A detailed summary of these datasets is presented in Table 1.
@@ 4 Methodology
In this section, we present the ValuesRAG which is specifically designed to address cultural biases and enhance contextual alignment in LLM-driven scenarios through a Retrieval-Augmented Generation (RAG) approach. ValuesRAG consists of three key components: Values and Demographic Summary Generation, which extracts and summarizes cultural values and demographic information from large-scale datasets (Section 4.1); Values-Augmented Generation, which incorporates these summaries into the generative process to align responses with the cultural context (Section 4.2); and Retrieval-based Values Alignment, which dynamically assigns relevant individual values to queries based on demographic profiles (Section 4.3). [^3] We provide a case study of retrieve in Appendix C. An overview of the ValuesRAG framework is provided in Figure 2.